OpenStudy (anonymous):

Find all solutions if 0° ≤ θ < 360°. When necessary, round your answers to the nearest tenth of a degree. (Enter your answers as a comma-separated list.) 5 cos^2 4θ + 5 sin 4θ = 4 I understand as much as 5(1-sin^24θ)+5sin4θ=4 5-5sin^24θ+5sin4θ=4 -5sin^24θ+5sin4θ+1=0 5sin^24θ-5sin4θ-1=0 quadratic formula: gets me 5(+/-) sqrt(45) /10 and after that i arcsin but it doesn't give me desired answers.

OpenStudy (dumbcow):

remember the input is 4x, they want solutions for "x" \[\sin(4x) = \frac{5 \pm \sqrt{45}}{10}\] \[x = \frac{\sin^{-1} (\frac{5 \pm \sqrt{45}}{10})}{4}\]

OpenStudy (dumbcow):

also arc sin yields 2 answers within domain [0,360] there will be total of 4 solutions

OpenStudy (dumbcow):

oops nevermind, only 2 the 5+sqrt45 is greater than 10 so arc sin of ratio greater than 1 is impossible

OpenStudy (anonymous):

I attached a file as an example of the problem of exactly where I got lost. In the video it stated -.366 somehow turned into 201.5 or 338.5 and quoted "because of sin" which was why I attempted arcsin but the results differ heavily.

OpenStudy (dumbcow):

oh ok http://www.wolframalpha.com/input/?i=arcsin%28-.366%29 this is result of arc sin --> look where it gives answer in degrees (-21.5) -21.5 = 360 -21.5 = 338.5 degrees